# Skagit Fisheries Enhancement Group   UMass Amherst

### v1

@@ Abstract
Object detectors often perform poorly on data that differs from their training set. Domain adaptive object detection (DAOD) methods have recently demonstrated strong results on addressing this challenge. Unfortunately, we identify systemic benchmarking pitfalls that call past results into question and hamper further progress: (a) Overestimation of performance due to underpowered baselines, (b) Inconsistent implementation practices preventing transparent comparisons of methods, and (c) Lack of generality due to outdated backbones and lack of diversity in benchmarks. We address these problems by introducing: (1) A unified benchmarking and implementation framework, Align and Distill (ALDI), enabling comparison of DAOD methods and supporting future development, (2) A fair and modern training and evaluation protocol for DAOD that addresses benchmarking pitfalls, (3) A new DAOD benchmark dataset, CFC-DAOD, enabling evaluation on diverse real-world data, and (4) A new method, ALDI++, that achieves state-of-the-art results by a large margin. ALDI++ outperforms the previous state-of-the-art by +3.5 AP50 on Cityscapes $\rightarrow$ Foggy Cityscapes, +5.7 AP50 on Sim10k $\rightarrow$ Cityscapes (where ours is the only method to outperform a fair baseline), and +2.0 AP50 on CFC Kenai $\rightarrow$ Channel. Our framework777github.com/justinkay/aldi, dataset888github.com/visipedia/caltech-fish-counting, and state-of-the-art method offer a critical reset for DAOD and provide a strong foundation for future research.
Keywords:
Domain adaptation Object detection
@@ 1 Introduction
Figure 1: ALDI achieves state-of-the-art performance in domain adaptive object detection (DAOD) and provides a unified framework for fair comparison. We show: (1) Inconsistent implementation practices give the appearance of steady progress in DAOD (left bars [12, 10, 8, 26, 34]); reimplementation and fair comparison with ALDI shows less difference between methods than previously reported (middle bars); (2) A fairly constructed source-only model (blue line) outperforms many existing DAOD methods, indicating less progress has been made than previously reported; and a proper oracle (orange line) outperforms all existing methods, in contrast to previously-published results; and (3) Our proposed method ALDI++ (green bars) achieves state-of-the-art performance on DAOD benchmarks such as Cityscapes normal-$\rightarrow$ Foggy Cityscapes and is complementary to ongoing advances in object detection like VitDet [33].
The challenge of DAOD. Modern object detector performance, though excellent across many benchmarks [36, 53, 52, 3, 49, 46], often severely degrades when test data exhibits a distribution shift with respect to training data [41]. For instance, detectors do not generalize well when deployed in new environments in environmental monitoring applications [30, 53]. Similarly, models in medical applications perform poorly when deployed in different hospitals or on different hardware than they were trained [55, 19]. Unfortunately, in real-world applications it is often difficult, expensive, or time-consuming to collect the additional annotations needed to address such distribution shifts in a supervised manner. An appealing option in these scenarios is unsupervised domain adaptive object detection (DAOD), which attempts to improve detection performance when moving from a “source” domain (used for training) to a “target” domain (used for testing) [32, 29] without the use of target-domain supervision.
@@
Impediments to progress. Recently-published results indicate DAOD is exceptionally effective, doubling the performance of source-only models and even outperforming fully-supervised oracles [34, 8, 5]. However, upon close examination we discover problems with current benchmarking practices that call these results into question:
@@
P2: Inconsistent implementation practices preventing transparent comparisons of methods. We find existing DAOD methods are built using a variety of different object detection libraries with inconsistent training settings, making it difficult to determine whether performance improvements come from new DAOD methods or simply improved hyperparameters. We find that tweaking these hyperparameters—whose values often differ between methods yet are not reported in papers—can lead to a larger change in performance than the proposed methods themselves (see Sec. 6.3), thus we cannot take reported advancements at face value. Without the ability to make fair comparisons we cannot transparently evaluate contributions nor make principled methodological progress.
P3: (a) Lack of diverse benchmarks and (b) outdated model architectures, leading to overestimation of methods’ generality. DAOD benchmarks have focused largely on urban driving scenarios with synthetic distribution shifts [48, 28], and methods continue to use outdated detector architectures for comparison with prior work [9]. The underlying assumption is that methods will perform equivalently across application domains and backbone architectures. We show that in fact the ranking of methods changes across benchmarks and architectures, revealing that published results may be uninformative for practitioners using modern architectures and real-world data.
@@
1. Align and Distill (ALDI), a unified benchmarking and implementation framework for DAOD. In order to enable fair comparisons, we first identify key themes in prior work (Sec. 2) and unify common components into a single state-of-the-art framework, ALDI (Sec. 3). ALDI facilitates detailed study of prior art and streamlined implementation of new methods, supporting future research.
2. A fair and modern training protocol for DAOD methods, enabled by ALDI. We provide quantitative evidence of the benchmarking pitfalls we identify and propose an updated training and evaluation protocol to address them (Sec. 6.1). This enables us to set more realistic and challenging targets for the DAOD community and perform the first fair comparison of prior work in DAOD (Sec. 6.2).
3. A new benchmark dataset, CFC-DAOD, sourced from a real-world adaptation challenge in environmental monitoring (Sec. 5). CFC-DAOD increases the diversity of DAOD benchmarks and is notably larger than existing options. We show that the ranking of methods changes across different benchmarks (Sec. 6.2), thus the community will benefit from an additional point of comparison.
@@ 2 Related Work
Two methodological themes have dominated recent DAOD research: feature alignment and self-training/self-distillation. We first give an overview of these themes and previous efforts to combine them, and then use commonalities to motivate our unified framework, Align and Distill, in Sec. 3.
@@
Feature alignment in DAOD. Feature alignment methods aim to make target-domain data “look like” source-domain data, reducing the magnitude of the distribution shift. The most common approach utilizes an adversarial learning objective to align the feature spaces of source and target data [16, 10, 9, 57]. Faster R-CNN in the Wild [9] utilizes adversarial networks at the image and instance level. SADA [10] extends this to multiple adversarial networks at different feature levels. Other approaches propose mining for discriminative regions [57], weighting local and global features differently [47], incorporating uncertainty [40], and using attention networks [51]. Alignment at the pixel level has also been proposed using image-to-image translation techniques to modify input images directly [12].
@@
Self-training/self-distillation in DAOD. Self-training methods use a “teacher” model to predict pseudo-labels on target-domain data that are then used as training targets for a “student” model. Self-training can be seen as a type of self-distillation [43, 6], which is a special case of knowledge distillation [25, 7] where the teacher and student models share the same architecture. Most recent self-training approaches in DAOD are based on the Mean Teacher [50] framework, in which the teacher model is updated as an exponential moving average (EMA) of the student model’s parameters. Extensions to Mean Teacher for DAOD include: MTOR, which utilizes graph structure to enforce student-teacher feature consistency [4], Probabilistic Teacher (PT), which uses probabilistic localization prediction and soft distillation losses [8], and Contrastive Mean Teacher (CMT), which uses MoCo [21] to enforce student-teacher feature consistency [5].
@@
Combining feature alignment and self-training. Several approaches utilize both feature alignment and self-training/self-distillation, motivating our unified framework. Unbiased Mean Teacher (UMT) [12] uses mean teacher in combination with image-to-image translation to align source and target data at the pixel level. Adaptive Teacher (AT) [55] uses mean teacher with an image-level discriminator network. Masked Image Consistency (MIC) [26] uses mean teacher, SADA, and a masking augmentation to enforce teacher-student consistency. Because these methods were implemented in different codebases using different training recipes and hyparameter settings, it is unclear which contributions are most effective and to what extent feature alignment and self-training are complementary. We address these issues by reimplementing these approaches in the ALDI framework and perform fair comparisons and ablation studies in Sec. 6.
@@
DAOD datasets. Cityscapes (CS) $\rightarrow$ Foggy Cityscapes (FCS) [11, 48] is a popular DAOD benchmark that emulates domain shift caused by changes in weather in urban driving scenarios. The dataset contains eight vehicle and person classes. Sim10k $\rightarrow$ CS [28] poses a Sim2Real challenge, adapting from video game imagery to real-world imagery. The benchmark focuses on a single class, “car”. Other common tasks include adapting from real imagery in PascalVOC [15] to clip art and watercolor imagery [27]. We report results on CS $\rightarrow$ FCS and Sim10k $\rightarrow$ CS due to their widespread popularity in the DAOD literature and focus on real applications. We note that existing benchmarks reflect a relatively narrow set of potential DAOD applications. To study whether methods generalize outside of urban driving scenarios, in Sec. 5 we introduce a novel dataset sourced from a real-world adaptation challenge in environmental monitoring, where imagery is much different from existing benchmarks.
@@ 3 Align and Distill (ALDI): Unifying DAOD
We first introduce Align and Distill (ALDI), a new benchmarking and implementation framework for DAOD. ALDI unifies existing approaches in a common framework, enabling fair comparisons and addressing P2. Inconsistent implementation practices, while also providing the foundation for development of a new method ALDI++ that achieves state-the-art performance (Sec. 4, Sec. 6.2). The framework is visualized in Fig. 2. All components are ablated in Sec. 6.3.
@@
Data. DAOD involves two datasets: a labeled source dataset $X_{src}$ and an unlabeled target dataset $X_{tgt}$. Each training step, a minibatch of size $B$ is constructed containing both $B_{src}$ source images and $B_{tgt}$ target images.
@@
Models. A student model $\theta_{stu}$ and a teacher model $\theta_{tch}$ are initialized with the same weights, typically obtained through supervised pretraining on ImageNet, COCO, or $X_{src}$. Pretraining on $X_{src}$ is often referred to as “burn-in.” The student is trained through backpropagation. The teacher’s weights are not updated by backpropagation, rather they are updated to be the EMA of student weights [50].
Training involves three objectives:
@@
1. Supervised training with source data. Each labeled source sample $x_{src,i}$ is transformed by some $t\sim T_{src}$, the set of possible source-domain transformations, then passed through the student model to obtain a supervised loss $L_{sup}$ given ground truth targets $y_{src,i}$. In our case $L_{sup}$ are Faster R-CNN losses [44].
2. Self-distillation with target data. Each unlabeled target sample $x_{tgt,i}$ passes through both the teacher and student models. The teacher’s predictions $\hat{p}_{tgt,i}$ act as distillation targets for the student’s predictions $p_{tgt,i}$, resulting in distillation losses $L_{distill}$ that are backpropagated through the student. Before computing $L_{distill}$ the teacher’s outputs are postprocessed to be either soft (e.g. logits or softmax outputs) or hard (e.g. thresholded pseudo-label) targets.
Before passing through the teacher, $x_{tgt,i}$ is transformed by some $\hat{t}\sim\hat{T}_{weak}$, a set of “weak” transformations—e.g. random horizontal flipping—that allow the teacher to provide high quality predictions $\hat{p}_{tgt}$. The same image is also passed to the student, this time transformed by $t\sim T_{tgt}$ which typically contains “stronger” augmentations such as color jitter or random erasing.
Table 1: ALDI unifies and extends existing work. We show settings to reproduce five prior works and our method ALDI++. See Fig. 2 and Sec. 3 for more details. Burn-in: fixed duration (Fixed), our approach (Ours, Sec. 4). Augs. $T_{src},T_{tgt}$: Random flip (F), multi-scale (M), crop & pad (CP), color jitter (J), cutout [13] (C), MIC [26]. ${\frac{1}{2}}$: augs used on half the images in the batch. $\frac{B_{tgt}}{B}$: Target-domain portion of minibatch of size $B$. Postprocess: Processing of teacher preds before distillation: sigmoid/softmax (Sharpen), sum class preds for pseudo-objectness (Sum), conf. thresholding (Thresh), NMS. $L_{distill}$: Distillation losses: hard pseudo-labels (Hard), continuous targets (Soft). $L_{align}$: Feature alignment losses: image-level adversarial (Img), instance-level adversarial (Inst), image-to-image translation (Img2Img). ${}^{\dagger}$: settings used in ALDI implementation (last column) but not in the original implementation (second-to-last column). ${}^{at}$: source-only and oracle results sourced from [34].
[…]
Burn-in $T_{src}$ $T_{tgt}$ $\frac{B_{tgt}}{B}$ process $L_{distill}$ $L_{align}$ (Reported) (w/ ALDI)
Source-only F, ${}^{\dagger}$, ${}^{\dagger}$, ${}^{\dagger}$ 0.0 ${}^{at}$23.5 59.1
SADA [10] F F 0.5 Img, Inst 44.0 54.2
PT [8] Fixed F F, J, C 0.3 Sharpen, Sum Soft 47.1 59.2
UMT [12] CP, J 0.5 Thresh, NMS Hard Img2Img 41.7 61.4
MIC [26] F F, J, MIC 0.5 Thresh, NMS Hard Img, Inst 47.6 61.7
[…]
ALDI++ Ours F, M, J, C F, M, J, MIC 0.5 Sharpen Soft 66.8
Oracle F, ${}^{\dagger}$, ${}^{\dagger}$, ${}^{\dagger}$ 1.0 ${}^{at}$42.7 67.2
@@
Figure 2: (Left) The ALDI student-teacher framework for DAOD. Each training step (moving left to right and bottom to top): (1) Sample $B_{src}$ labeled source images $x_{src}$; transform by $t\sim T_{src}$; pass to student; compute supervised loss $L_{sup}$ using ground-truth labels $y_{src}$. (2) Sample $B_{tgt}$ unlabeled target images $x_{tgt}$; transform by $t\sim T_{tgt}$; pass to student to get preds $p_{tgt}$. Compute alignment objectives $L_{align}$ using $x_{src}$ and $x_{tgt}$. (3) Pass same unlabeled target data $x_{tgt}$, weakly transformed, to teacher; postprocess to obtain teacher preds $\hat{p}_{tgt}$. Compute distillation loss $L_{distill}$ between teacher and student preds. Use stop gradient (SG) on teacher model; update teacher to the EMA of student’s weights. (Middle, Right) ALDI++ (Sec. 4) introduces two new methods to achieve state-of-the-art performance: (Middle) A robust burn-in strategy utilizing strong augmentations and EMA, and (Right) Multi-task soft distillation losses to train the student using teacher outputs at all detector stages. $\sigma$: sigmoid or softmax for binary cross-entropy and cross entropy losses, respectively.
@@
3. Feature alignment. $x_{src,:}$ and $x_{tgt,:}$ are “aligned” via an alignment objective $L_{align}$ that enforces invariance across domains either at the image or feature level.
@@
Unification of prior work. We demonstrate the generality of our framework by reimplementing five recently-proposed methods on top of ALDI for fair comparison: UMT [12], SADA [10], PT [8], MIC [26], and AT [34]. In Tab. 1 we enumerate the settings required to reproduce each method.

### v3

# Align and Distill: Unifying and Improving Domain Adaptive Object Detection

@@ Abstract
Object detectors often perform poorly on data that differs from their training set. Domain adaptive object detection (DAOD) methods have recently demonstrated strong results on addressing this challenge. Unfortunately, we identify systemic benchmarking pitfalls that call past results into question and hamper further progress: (a) Overestimation of performance due to underpowered baselines, (b) Inconsistent implementation practices preventing transparent comparisons of methods, and (c) Lack of generality due to outdated backbones and lack of diversity in benchmarks. We address these problems by introducing: (1) A unified benchmarking and implementation framework, Align and Distill (ALDI), enabling comparison of DAOD methods and supporting future development, (2) A fair and modern training and evaluation protocol for DAOD that addresses benchmarking pitfalls, (3) A new DAOD benchmark dataset, CFC-DAOD, increasing the diversity of available DAOD benchmarks, and (4) A new method, ALDI++, that achieves state-of-the-art results by a large margin. ALDI++ outperforms the previous state-of-the-art by +3.5 AP50 on Cityscapes $\rightarrow$ Foggy Cityscapes, +5.7 AP50 on Sim10k $\rightarrow$ Cityscapes (where ours is the only method to outperform a fair baseline), and +0.6 AP50 on CFC-DAOD. ALDI and ALDI++ are architecture-agnostic, setting a new state-of-the-art for YOLO and DETR-based DAOD as well without additional hyperparameter tuning. Our framework222github.com/justinkay/aldi, dataset333github.com/visipedia/caltech-fish-counting, and method offer a critical reset for DAOD and provide a strong foundation for future research.
@@ 1 Introduction
The challenge of DAOD. Modern object detector performance, though excellent across many benchmarks (Lin et al., 2014; Weinstein et al., 2021b; a; Bondi et al., 2018; Schneider et al., 2018; Rodriguez et al., 2011), often severely degrades when test data exhibits a distribution shift with respect to training data (Oza et al., 2023). For instance, detectors do not generalize well when deployed in new environments in environmental monitoring applications (Kay et al., 2022; Weinstein et al., 2021b). Similarly, models in medical applications perform poorly when deployed in different hospitals or on different hardware than they were trained (Xue et al., 2023; Guan & Liu, 2021). Unfortunately, in real-world applications it is often difficult, expensive, or time-consuming to collect the additional annotations needed to address such distribution shifts in a supervised manner. An appealing option in these scenarios is unsupervised domain adaptive object detection (DAOD), which attempts to improve detection performance when moving from a “source” domain (used for training) to a “target” domain (used for testing) (Koh et al., 2021; Kalluri et al., 2023) without the use of target-domain supervision.
@@
Impediments to progress. Recently-published results indicate DAOD is exceptionally effective, doubling the performance of source-only models and even outperforming fully-supervised oracles (Li et al., 2022b; Chen et al., 2022; Cao et al., 2023). However, upon close examination we discover problems with current benchmarking practices that call these results into question:
@@
P2: Inconsistent implementation practices preventing transparent comparisons of methods. We find existing DAOD methods are built using a variety of different object detection libraries with inconsistent training settings, making it difficult to determine whether performance improvements come from new DAOD methods or simply improved hyperparameters. We find that tweaking these hyperparameters—whose values often differ between methods yet are not reported in papers—can lead to a larger change in performance than the proposed methods themselves (see Section 6.3), thus we cannot take reported advancements at face value. Without the ability to make fair comparisons we cannot transparently evaluate contributions nor make principled methodological progress.
P3: (a) Lack of diverse benchmarks and (b) outdated backbone architectures, leading to overestimation of methods’ generality. DAOD benchmarks have focused largely on urban driving scenarios with synthetic distribution shifts (Sakaridis et al., 2018; Johnson-Roberson et al., 2016), and methods continue to use outdated detector backbones for comparison with prior work (Chen et al., 2018). We show that in fact the ranking of methods changes across benchmarks and architectures, revealing that published results may be uninformative for practitioners using modern architectures and real-world data.
Figure 1: ALDI provides a unified framework for fair comparison of domain adaptive object detection (DAOD) methods, and ALDI++ achieves state-of-the-art performance. We show: (1) Inconsistent implementation practices give the appearance of steady progress in DAOD (left bars (Deng et al., 2021; Chen et al., 2021; 2022; Hoyer et al., 2023; Li et al., 2022b)); reimplementation and fair comparison with ALDI shows less difference between methods than previously reported (middle bars); (2) A fairly constructed source-only model (blue line) outperforms many existing DAOD methods, indicating less progress has been made than previously reported; and a proper oracle (orange line) outperforms all existing methods, in contrast to previously-published results; and (3) Our proposed method ALDI++ (green bars) achieves state-of-the-art performance on DAOD benchmarks such as Cityscapes $\rightarrow$ Foggy Cityscapes and is complementary to ongoing advances in object detection like VitDet (Li et al., 2022a).
@@
1. Align and Distill (ALDI), a unified benchmarking and implementation framework for DAOD. In order to enable fair comparisons, we first identify key themes in prior work (Section 2) and unify common components into a single state-of-the-art framework, ALDI (Section 3). ALDI facilitates detailed study of prior art and streamlined implementation of new methods, supporting future research.
2. A fair and modern training protocol for DAOD methods, enabled by ALDI. We provide quantitative evidence of the benchmarking pitfalls we identify and propose an updated training and evaluation protocol to address them (Section 6.1). This enables us to set more realistic and challenging targets for the DAOD community and perform the first fair comparison of prior work in DAOD (Section 6.2).
3. A new benchmark dataset, CFC-DAOD, sourced from a real-world adaptation challenge in environmental monitoring (Section 5). CFC-DAOD increases the diversity of DAOD benchmarks and is notably larger than existing options. We show that the ranking of methods changes across different benchmarks (Section 6.2), thus the community will benefit from an additional point of comparison.
@@ 2 Related Work
Our work concerns domain adaptive 2D object detection (DAOD). Two methodological themes have dominated recent DAOD research: feature alignment and self-training/self-distillation. We first give an overview of these themes and previous efforts to combine them, and then use commonalities to motivate our unified framework, Align and Distill, in Section 3.
@@
Feature alignment in DAOD. Feature alignment methods aim to make target-domain data “look like” source-domain data, reducing the magnitude of the distribution shift. The most common approach utilizes an adversarial learning objective to align the feature spaces of source and target data (Ganin & Lempitsky, 2015; Chen et al., 2021; 2018; Zhu et al., 2019). Faster R-CNN in the Wild (Chen et al., 2018) utilizes adversarial networks at the image and instance level. SADA (Chen et al., 2021) extends this to multiple adversarial networks at different feature levels. Other approaches propose mining for discriminative regions (Zhu et al., 2019), weighting local and global features differently (Saito et al., 2019), incorporating uncertainty (Nguyen et al., 2020), and using attention networks (Vs et al., 2021). Alignment at the pixel level has also been proposed using image-to-image translation techniques to modify input images directly (Deng et al., 2021).
@@
Self-training/self-distillation in DAOD. Self-training methods use a “teacher” model to predict pseudo-labels on target-domain data that are then used as training targets for a “student” model. Self-training can be seen as a type of self-distillation (Pham et al., 2022; Caron et al., 2021), which is a special case of knowledge distillation (Hinton et al., 2015; Chen et al., 2017) where the teacher and student models share the same architecture. Most recent self-training approaches in DAOD are based on the Mean Teacher (Tarvainen & Valpola, 2017) framework, in which the teacher model is updated as an exponential moving average (EMA) of the student model’s parameters. Extensions to Mean Teacher for DAOD include: MTOR, which utilizes graph structure to enforce student-teacher feature consistency (Cai et al., 2019), Probabilistic Teacher (PT), which uses probabilistic localization prediction and soft distillation losses (Chen et al., 2022), and Contrastive Mean Teacher (CMT), which uses MoCo (He et al., 2020) for student-teacher consistency (Cao et al., 2023).
@@
Combining feature alignment and self-training. Several approaches utilize both feature alignment and self-training/self-distillation, motivating our unified framework. Unbiased Mean Teacher (UMT) (Deng et al., 2021) uses mean teacher in combination with image-to-image translation to align source and target data at the pixel level. Adaptive Teacher (AT) (Xue et al., 2023) uses mean teacher with an image-level discriminator network. Masked Image Consistency (MIC) (Hoyer et al., 2023) uses mean teacher, SADA, and a masking augmentation to enforce teacher-student consistency. Because these methods were implemented in different codebases using different training recipes and hyperparameter settings, it is unclear which contributions are most effective and to what extent feature alignment and self-training are complementary. We address these issues by reimplementing these approaches in the ALDI framework and perform fair comparisons and ablation studies in Section 6.
DAOD implementations. There are two components to an object detector design: the detection architecture (e.g. Faster R-CNN Ren et al. (2015), YOLO Redmon et al. (2016), DETR Carion et al. (2020)) and the backbone (e.g. VGG Simonyan & Zisserman (2014), ResNet He et al. (2016), ViT Dosovitskiy et al. (2020)). Current state-of-the-art methods in DAOD predominantly use Faster R-CNN architectures. DOAD methods for YOLO and DETR backbones have recently received some attention Zhou et al. (2023); Yu et al. (2022); Jia et al. (2023), but have yet to surpass Faster R-CNN-based methods’ performance. For this reason, our main experiments also utilize the Faster R-CNN architecture. Existing methods differ in their choice of backbone, making comparisons difficult; we address this by consistently utilizing ResNet-50 backbones for all experiments and in our re-implementations of prior work. However, the ALDI framework is architecture and backbone agnostic, and we provide additional experiments using YOLO and DETR architectures, as well as ViT and ConvNeXt Liu et al. (2022) backbones.
@@
DAOD datasets. Cityscapes (CS) $\rightarrow$ Foggy Cityscapes (FCS) (Cordts et al., 2016; Sakaridis et al., 2018) is a popular DAOD benchmark that emulates domain shift caused by changes in weather in urban driving scenarios. The dataset contains eight vehicle and person classes. Sim10k $\rightarrow$ CS (Johnson-Roberson et al., 2016) poses a Sim2Real challenge, adapting from video game imagery to real-world imagery. The benchmark focuses on a single class, “car”. Other common tasks include adapting from real imagery in PascalVOC (Everingham et al., 2010) to clip art and watercolor imagery (Inoue et al., 2018). We report results on CS $\rightarrow$ FCS and Sim10k $\rightarrow$ CS due to their widespread popularity in the DAOD literature and focus on real applications. We note that existing benchmarks reflect a relatively narrow set of potential DAOD applications. To study whether methods generalize outside of urban driving scenarios, in Section 5 we introduce a novel dataset sourced from a real-world adaptation challenge in environmental monitoring, where imagery is much different from existing benchmarks.
@@ 3 Align and Distill (ALDI): Unifying DAOD
We first introduce Align and Distill (ALDI), a new benchmarking and implementation framework for DAOD. ALDI unifies feature alignment and self-distillation approaches in a common framework, enabling fair comparisons and addressing P2. Inconsistent implementation practices, while also providing the foundation for development of a new method ALDI++ that achieves state-the-art performance (Section 4, Section 6.2). The framework is visualized in Fig. 2. All components are ablated in Section 6.3.
@@
Data. DAOD involves two datasets: a labeled source dataset $X_{\text{src}}$ and an unlabeled target dataset $X_{\text{tgt}}$. Each training step, a minibatch of size $B$ is constructed containing both $B_{\text{\text{src}}}$ source images and $B_{\text{\text{tgt}}}$ target images, $B=B_{\text{src}}+B_{\text{tgt}}$.
@@
Models. ALDI is designed as a student-teacher framework to facilitate algorithms utilizing self-training/self-distillation. When enabled, both a student model $\theta_{\text{stu}}$ and a teacher model $\theta_{\text{tch}}$ are initialized with the same weights, typically obtained through supervised pretraining on ImageNet or $X_{\text{src}}$. Pretraining on $X_{\text{src}}$ is often referred to as “burn-in.” The student is trained via backpropagation, while the teacher’s weights are updated each training step to be the EMA of the student’s weights (Tarvainen & Valpola, 2017), i.e. $\theta_{\text{tch}}=\alpha\theta_{\text{tch}}+(1-\alpha)\theta_{\text{stu}}$ with $\alpha\in[0,1]$. After training, we keep $\theta_{\text{tch}}$ and discard $\theta_{\text{stu}}$. Algorithms that do not use self-training/self-distillation (e.g. SADA (Chen et al., 2021)) simply disable $\theta_{\text{tch}}$.
In this paper we focus predominantly on two-stage detectors based on Faster R-CNN (Ren et al., 2015) as they are currently the state-of-the-art in DAOD, though we note that our framework is architecture-agnostic and also supports YOLO and DETR-based detectors. We provide additional YOLO and DETR results in Section A.1.
Training involves one or more of the following three objectives. We note that each objective is optional in order to support a range of algorithmic approaches.
@@
1. Supervised training with source data. For each labeled source sample $x_{\text{src},i}$, we apply a transformation $t\sim T_{\text{src}}$, where $T_{\text{src}}$ is the set of possible source-domain transformations. The transformed sample is passed through the student model to compute the supervised loss $\mathcal{L}_{\text{sup}}$ given the ground truth targets $y_{\text{src},i}$:
$\mathcal{L}_{\text{sup}}=\frac{1}{B_{\text{src}}}\sum_{i=1}^{B_{\text{src}}}\mathcal{L}\left(\theta_{\text{stu}}(t(x_{\text{src},i})),y_{\text{src},i}\right)$ (1)
where $\mathcal{L}(\cdot,\cdot)$ are standard object detection loss functions, e.g. those of Faster R-CNN (Ren et al., 2015).
2. Self-distillation with target data. For each unlabeled target sample $x_{\text{tgt},i}$, we transform the input using $\hat{t}\sim T_{\text{weak}}$ (a set of weak transformations) for the teacher model and $t\sim T_{\text{tgt}}$ (stronger transformations) for the student model. The teacher’s predictions $\hat{p}_{\text{tgt},i}$ serve as distillation targets for the student’s predictions $p_{\text{tgt},i}$, and we compute distillation loss $\mathcal{L}_{\text{distill}}$ :
$\hat{p}_{\text{tgt},i}=\theta_{\text{tch}}(\hat{t}(x_{\text{tgt},i}))$ (2)
$p_{\text{tgt},i}=\theta_{\text{stu}}(t(x_{\text{tgt},i}))$ (3)
$\mathcal{L}_{\text{distill}}=\frac{1}{B_{\text{tgt}}}\sum_{i=1}^{B_{\text{tgt}}}\mathcal{L}_{\text{distill}}\left(p_{\text{tgt},i},\hat{p}_{\text{tgt},i}\right)$ (4)
where teacher outputs $\hat{p}_{\text{tgt},i}$ are postprocessed to be either soft (e.g., logits or softmax outputs) or hard (e.g., thresholded pseudo-label) targets and the choice of distillation loss is method-specific. This formulation unifies different distillation techniques into a common objective, supporting a range of approaches.
ALDI module Example settings
Network init. ImageNet (Chen et al., 2021; Deng et al., 2021; Hoyer et al., 2023),
$\theta_{stu},\theta_{tch}$ Source-domain burn-in (Chen et al., 2022; Li et al., 2022b)
$T_{src}$, Flip and scale, SimCLR augs (Chen et al., 2022; Li et al., 2022b),
$T_{tgt}$ Image masking (Hoyer et al., 2023)
$1:1$ (Chen et al., 2021; Deng et al., 2021; Hoyer et al., 2023),
$B_{src}:B_{tgt}$ $2:1$ (Chen et al., 2022; Li et al., 2022b)
Pseudo-label (Deng et al., 2021; Hoyer et al., 2023; Li et al., 2022b),
$L_{distill}$ Knowledge distillation (Chen et al., 2022)
Adversarial (Li et al., 2022b; Chen et al., 2021; Hoyer et al., 2023),
$L_{align}$ Image-to-image translation (Deng et al., 2021)
@@
Figure 2: (Left) The ALDI framework. Each training step (moving left to right and bottom to top): (1) Sample $B_{\text{src}}$ labeled source images $x_{\text{src}}$; transform by $t\sim T_{\text{src}}$; pass to student; compute supervised loss $L_{\text{sup}}$ using ground-truth labels $y_{\text{src}}$. (2) Sample $B_{\text{tgt}}$ unlabeled target images $x_{\text{tgt}}$; transform by $t\sim T_{\text{tgt}}$; pass to student to get preds $p_{\text{tgt}}$. Compute alignment objectives $L_{\text{align}}$ using $x_{\text{src}}$ and $x_{\text{tgt}}$. (3) Pass same unlabeled target data $x_{\text{tgt}}$, weakly transformed, to teacher; postprocess to obtain teacher predictions $\hat{p}_{\text{tgt}}$. Compute distillation loss $L_{\text{distill}}$ between teacher and student predictions. Use stop gradient (SG) on teacher model; update teacher to the EMA of student’s weights. (Right) Example settings for each component of ALDI. ALDI supports a range of existing methods off-the-shelf while providing a general implementation framework for new methods.
@@
3. Feature alignment. The source samples $x_{\text{src},i}$ and target samples $x_{\text{tgt},i}$ are optionally “aligned” using an alignment objective $\mathcal{L}_{\text{align}}$ that enforces invariance across domains at either the image or feature level. This formulation is general; however, in this paper, we focus on two common alignment losses: domain-adversarial training and image-to-image alignment.
Domain-adversarial training (i.e. DANN (Ganin & Lempitsky, 2015)) trains a domain classifier $D$ to distinguish between source and target features, while the feature extractor aims to confuse $D$:
$\mathcal{L}_{\text{align,DANN}}=-\frac{1}{B}\sum_{i=1}^{B}\left[y_{\text{dom},i}\log(D({\theta}(x_{i})))+(1-y_{\text{dom},i})\log(1-D({\theta}(x_{i})))\right]$ (5)
where $y_{\text{dom},i}$ is the domain label (source = 0, target = 1) and ${\theta}(x_{i})$ is a feature representation of $x_{i}$.
Image-to-image alignment instead pursues domain invariance in the pixel space. Given image-to-image generative models $G_{\text{src}}$, $G_{\text{tgt}}$ (e.g., a CycleGAN (Zhu et al., 2017)), images are “translated” (a pixel-level transformation) from the source domain to the target domain and vice versa. We then obtain $x_{\text{tgt-like},i}=G_{\text{src}}\left(x_{\text{src},i}\right)$, $x_{\text{src-like},i}=G_{\text{tgt}}\left(x_{\text{tgt},i}\right)$, and substitute into Eq. 1, Eq. 2, and Eq. 3.
@@
Unification of prior work. We demonstrate the generality of our framework by reimplementing five recently-proposed methods on top of ALDI for fair comparison: UMT (Deng et al., 2021), SADA (Chen et al., 2021), PT (Chen et al., 2022), MIC (Hoyer et al., 2023), and AT (Li et al., 2022b). We enumerate the settings required to reproduce each method in Appendix C.